OpenStudy (anonymous):

Write the missing factor. (d - 5)( ? ) = d^2- 25 A. 5 - d B. 5 + 5d C. d - 5 D. d + 5

OpenStudy (anonymous):

Um we know the following: \( (a+b)(a-b)=a^2-b^2\) In your particular case you have \(d^2-25=d^2-5^2\) Using the above information we have \[ (d+5) ( ?)=d^2-5^2\) Can you take a guess what belongs in the bracket?

OpenStudy (anonymous):

d + 5?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

Can you help me with one more? @samjordon

OpenStudy (anonymous):

waitttttt

OpenStudy (anonymous):

that was wrong ,,,,,

OpenStudy (anonymous):

look at my example (a+b)(a-b)=a^2+b^2 the signs between the 2 brackets must differ

OpenStudy (anonymous):

d - 5?

OpenStudy (anonymous):

Yes sorry

OpenStudy (anonymous):

thats the correct answer

OpenStudy (anonymous):

When factoring a trinomial of the form ax2 + bx + c where a, b, and c are all positive, you need to find two numbers, m and n, such that ax2 + bx + c = ax2 + mx + nx + c and __________. A.m and n are both positive B.m and n are both negative C.m is positive, n is negative, and m > n D.m is positive, n is negative, and n > m

OpenStudy (anonymous):

I would most likely think its A Since a and c is positive and m*n=a*c then m and n both must be negative or both must positive Since b is positive and m+n=b so we can conclude that m and n must be positive